OpenStudy (anonymous):

Solve for x. 3 over x equals x over 32 96 48 4 radical 6 6 radical 4

OpenStudy (zzr0ck3r):

\[\frac{3}{x}=\frac{x}{32}\] multiply both sides by 32x and tell me what you have

OpenStudy (tpaulus):

Setup a proportion, then cross-multiply and divide

OpenStudy (anonymous):

96

OpenStudy (zzr0ck3r):

what about the other side?

OpenStudy (zzr0ck3r):

\[32x\frac{3}{x}=32x\frac{x}{32}\rightarrow 96=\frac{32}{32}*x^2=?\]

OpenStudy (zzr0ck3r):

\[96=32x\frac{x}{32}\] just like the x's cancel out on the left, the 32's cancel out on the right \[96=x^2\]

OpenStudy (anonymous):

Ohh okay, I kind of understand now.

OpenStudy (zzr0ck3r):

\[\pm 4\sqrt{6}\]

OpenStudy (anonymous):

So I was right ? :D

OpenStudy (zzr0ck3r):

plus or minus 4 radical 6

OpenStudy (zzr0ck3r):

I think you just said 4 radical 6

OpenStudy (anonymous):

Yeah o.O